OpenStudy (kayders1997):

What is the derivative of (arctan(3x))/x

zepdrix (zepdrix):

Hey Kay :) Do you have the derivative of arctan(x) handy? We'll need that at some point.

zepdrix (zepdrix):

Let's get our derivative "set up" first though, and we'll worry about the arctangent afterwards.

OpenStudy (kayders1997):

The derivative of arctan(x) is 1/1x^2 if I'm correct

zepdrix (zepdrix):

1/(1+x^2) I think you made a silly typo there, ya? :)

OpenStudy (kayders1997):

Yeah

zepdrix (zepdrix):

So we have a quotient, we'll apply quotient rule, ya?\[\large\rm \frac{d}{dx}\frac{\arctan(3x)}{x}=\frac{\color{royalblue}{\left[\arctan(3x)\right]'}x-\arctan(3x)\color{royalblue}{\left[x\right]'}}{(x)^2}\]Understand the set up? We have to differentiate the blue stuff.

OpenStudy (kayders1997):

I don't know if I'm supposed to do something... but I can't see it all?

zepdrix (zepdrix):

You mean you can't see the fancy math formatted? Hmm are you using Internet Explorer? That browser doesn't load the plugin correctly sometimes :(

OpenStudy (kayders1997):

I'm just using internet on my samsung, thats all it says

zepdrix (zepdrix):

I can use the draw tool instead if it's not showing up properly

OpenStudy (kayders1997):

Would using a computer maybe wrk better?

zepdrix (zepdrix):

Ugh, website is freezing again :(

OpenStudy (kayders1997):

some of it

zepdrix (zepdrix):

So we're using our quotient rule.